OpenStudy (anonymous):

1. Arrange the list of numbers in increasing order 4/5 7/9 6/7 A. 7/9 4/5 6/7 <----- B. 4/5 7/9 6/7 C.4/5 6/7 7/9 D.6/7 4/5 7/9

OpenStudy (anonymous):

@satellite73 @iambatman @acxbox22

OpenStudy (anonymous):

4/5 = 0.8 7/9 = 0.77... 6/7 = .856

OpenStudy (anonymous):

that looks good

OpenStudy (anonymous):

Yeah now it's correct, as satellite pointed out :p

OpenStudy (anonymous):

So B? :)

OpenStudy (anonymous):

no dear, A like you had

OpenStudy (anonymous):

...You said I was wrong.....

OpenStudy (anonymous):

That's because on your previous post, you had it at C or something else...

OpenStudy (anonymous):

A is the correct answer the last time you had a different one i think don't fret, moveon.org

OpenStudy (anonymous):

Er...No I didn't. I just copied and pasted it here. -_-

OpenStudy (anonymous):

ok then i made a mistake it is A

OpenStudy (anonymous):

New question

OpenStudy (anonymous):

KK, do u want me to post it on as a new question or here?

OpenStudy (anonymous):

new thread spend the site